OpenStudy (baby456):

adding and subtracting polynomials! 39.(-3x^2+4x+7)+(7x^2-3x-12) 40.(2x^2-5x-9)-(43x^2-7x+12) 41.(x^2-4x-11)+(8x^2+11x+15) 42.(4x^2-9x-12)-(2x^2+9x-11)

OpenStudy (baby456):

Please help me!

OpenStudy (anonymous):

39. \((-3x^{2}+7x^{2}) + (4x-3x) +(7-12) = ?\)

OpenStudy (baby456):

@satellite73

OpenStudy (anonymous):

39. -3x^2+4x+7+7x^2-2x-12 = -3x^2+7x^2+4x-2x+7-12 = 4x^2+2x-5 If you understand what I did there, all you need to do is follow the same kind of principles then you should get the rest. With 40. and 42. all you need to do is picture that a -1 is in front of the second bracket and multiply that -1 by everything inside the bracket then follow the same pattern as I have shown previously.

OpenStudy (anonymous):

Just make sure to collect like terms, which means you have to add/subtract all the terms that have a x^2, x, and a base number together.

OpenStudy (anonymous):

39. 4x^2+x-5 40. -41x^2-12x+3 41. 9x^2+7x+4 42. 2x^2-23

OpenStudy (baby456):

thanks guys

OpenStudy (baby456):

but can you explain how you got 40.

OpenStudy (anonymous):

combine light terms

OpenStudy (anonymous):

2x^2+-43x^2=-41x^2 -5x+-7x=-12x -9+12=3

OpenStudy (anonymous):

(2x^2-5x-9)-(43x^2-7x+12) = 2x^2-5x-9-43x^2+7x-12 = 2x^2-43x^2-5x+7x-9-12 = 41x^2+2x-21 You have to make sure to get all the like terms together. Also make sure to distribute the negative into the brackets which will affect everything else inside the brackets.

OpenStudy (anonymous):

Sorry, -41x^2+2x-21.

OpenStudy (baby456):

and 42.

OpenStudy (baby456):

well i kind of get it now thank you!

OpenStudy (anonymous):

shonari ur wrong im right

OpenStudy (anonymous):

You forgot to distribute the negative into the brackets, therefore giving you a positive 7 rather than a negative 7.

OpenStudy (baby456):

i guess i will show it to my teacher tommorow to see which one is right!

OpenStudy (anonymous):

Hopefully all goes well, good luck.

OpenStudy (anonymous):

when u combine -5-7 into the calculator it will give u -12 not 2

OpenStudy (baby456):

i could check it online!

OpenStudy (anonymous):

-(43x^2-7x+12) means that you will multiply -1 and 43, -1 and -7 and -1 and 12.

OpenStudy (baby456):

2x2−5x−9−(43x2−7x+12) =2x2+(−5x)+(−9)+(−43x2)+7x+(−12) =(−41x2)+2x+(−21) =−41x2+2x−21 source @ http://www.mathpapa.com/algebra-calculator.html

OpenStudy (anonymous):

sorry ur right i didn't c the (-)

OpenStudy (baby456):

so @shonari_d was right!

OpenStudy (anonymous):

It's alright survivor, we all make mistakes!

OpenStudy (baby456):

if i could i would give survivor a medal too!

OpenStudy (baby456):

thank you guys so much! i understand more than i did a ouple of dys ago! bye!

OpenStudy (anonymous):

Anytime, cya later!

OpenStudy (anonymous):

http://www.acalculator.com/math-calculators.html